OpenStudy (anonymous):

Simplify the expression. If the simplified expression is written in standard form, what is the leading coefficient?

OpenStudy (anonymous):

Firstly solve for : \[\large x \sqrt{64x^3} \implies x \sqrt{8 \times 8 \times x \times x \times x}\]

OpenStudy (anonymous):

So you can take one 8 and one x out: \[\large \implies x \times 8 \times x \sqrt{x}\] \[\large \implies 8x^2 \sqrt{x}\]

OpenStudy (anonymous):

So the leading coefficient is 8?

OpenStudy (anonymous):

This is for first term only..

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

No need to solve for second term as it is simplified.. Now for third you can write: \[\large 8 \sqrt{2x} \implies 8 \sqrt{2} \times \sqrt{x} \implies 8 \sqrt{2} \cdot \sqrt{x}\]

OpenStudy (anonymous):

Now you can take root(x) as common from all: \[\implies \color{green}{\large \sqrt{x}[8x^2 + 2x + 8\sqrt{2}]}\]

OpenStudy (anonymous):

In my view It cannot be factorized further..

OpenStudy (anonymous):

Okay. But the leading coefficient is still 8 right?

OpenStudy (anonymous):

Do you have answer choices with you??

OpenStudy (anonymous):

no its a fill in the blank

OpenStudy (anonymous):

I am guessing why are you concentrating on 8 ?

OpenStudy (anonymous):

because its the first coefficient

OpenStudy (anonymous):

yes its 8

OpenStudy (anonymous):

okay thank you!